How much agarose would you need to prepare a 30 mL, 1.5% (w/v) agarose gel?
Answer #1

The percentage measurement is a weight/volume thing . For example, a 1% gel would be 1 g agarose in 100 ml TAE ( TAE is Tris Acetate EDTA, it is a buffer and we make gels with TAE and run them in TAE buffer).

1 % means 1 g in 100 ml

Let x gram of agarose in 30 ml gives 1.5 % agarose gel

ie, x / 30 = 1.5 %

x / 30 = 1.5 / 100

Therefore x = ( 1.5 X 30 ) ÷ 100

x = 0.45 gram

0.45 g of agarose is needed to a 30ml, 1.5 % agarose gel.